Join the Boy Scouts for a FREE pancake breakfast

On Saturday, September 2 the Boy Scouts of Troop 171 will have a pancake breakfast fundraiser to benefit the Aurora Public Library. Coffee, juice, pancakes and eggs will be served. The cost is free, but in order to help the library we are accepting donations of books and money. Meet us in the parking lot of Trinity Baptist Church at 15555 E. Quincy Avenue from 7:00–11:00. A splendid time is guaranteed for all.

Please join us for good food and good fellowship. Your presence not only helps the library, it helps a scout to achieve the rank of Eagle. Direct any questions to (303) 766–3525.